### v3

# Resolving power: A general approach to compare the distinguishing ability of threshold-free evaluation metrics

Abstract
Selecting an evaluation metric is fundamental to model development, but uncertainty remains about when certain metrics are preferable and why. This paper introduces the concept of resolving power to describe the ability of an evaluation metric to distinguish between binary classifiers of similar quality. This ability depends on two attributes: 1. The metric’s response to improvements in classifier quality (its signal), and 2. The metric’s sampling variability (its noise). The paper defines resolving power generically as a metric’s sampling uncertainty scaled by its signal. A simulation study compares the area under the receiver operating characteristic curve (AUROC) and the and the area under the precision-recall curve (AUPRC) in a variety of contexts. It finds that the AUROC generally has greater resolving power, but that the AUPRC is better when searching among high-quality classifiers applied to low prevalence outcomes. The paper also proposes an empirical method to estimate resolving power that can be applied to any dataset and any initial classification model. The AUROC is useful for developing the resolving power concept, but it has been criticized for being misleading. Newer metrics developed to address its interpretative issues can be easily incorporated into the resolving power framework. The best metrics for model search will be both interpretable and high in resolving power. Sometimes these objectives will conflict and how to address this tension remains an open question.
keywords:
Evaluation metrics, Binary classification, Receiver Operating Characteristic, Precision-Recall
1 Introduction
There is a large and growing collection of evaluation metrics used for binary classification models. Choosing a metric can be challenging as model evaluation serves a variety of goals. One is interpretation, meaning that the metric is both easily understood and sensitive to aspects of quality that are relevant to the user. Simple classification accuracy, for example, is misleading for outcomes that are both low prevalence and relatively severe, such as rare but serious diseases. Evaluation metrics are also used to select the best model from a collection of competitors [15]. This includes selection between different model classes, such as between a simple baseline model and more complex machine learning models. And it includes selection within a model class, as occurs with hyperparameter search during model tuning. Another goal of evaluation is to estimate how well a given model will perform on future, unseen cases [17].
The Receiver Operating Characteristic (ROC) curve has become a favored method for evaluating binary classification models, in part due to the shortcomings of simple classification accuracy [6]. More recently, many have argued that the precision-recall curve (PRC) is preferable when there is a strong class imbalance and where there is low value in true-negative predictions [1, 17, 4]. Relative to the ROC curve, the PRC gives more weight to the highest-ranked cases located in the “early retrieval area” of ROC space. These cases are especially important when capacity to act is limited, such as when a health system has resources to intervene on only their sickest patients.
Sampling uncertainty is a neglected aspect of model evaluation within the field of machine learning [19], but it is essential to account for when data is limited [1, 5]. There has been scant research that compares the sampling precision of evaluation metrics for binary classifiers. One exception is found in Zhou [21], who used a link prediction task on a toy network model with a tunable noise parameter. For these network models Zhou finds that the area under the ROC curve (AUROC) and the area under the PRC (AUPRC) are much more discriminating than “balanced precision”, and that the AUROC is slightly more discriminating than AUPRC. [^1] Zhou finds balanced precision by choosing the decision threshold so that precision equals recall.
This paper pursues a general approach for comparing evaluation metrics. It also seeks specific conclusions about when and by how much some metrics are better than others. This project is conceptually difficult since evaluation metrics themselves are used to measure quality, each encoding different assumptions about what makes a model better or worse. The paper’s strategy is to use a collection of sampling models to construct a quality dimension that serves as the common standard of comparison. The sampling models are used to assess how an evaluation metric responds to changes in model quality (its signal) and how much variability it has at a given level of quality (its noise). These two quantities are combined to form an evaluation metric’s resolving power, which is a type of signal-to-noise ratio. The resolving power of a microscope is its capacity to distinguish between two close objects. By analogy, the resolving power of an evaluation metric describes how well it differentiates between models of similar quality. More specifically, resolving power is defined as a metric’s sampling uncertainty mapped to a common scale.
Resolving power draws inspiration from several previous works such as Saito and Rehmsmeier [17], Mazzanti [14], and hand2023notes. Each of these analyses addresses a metric’s adequacy in describing performance. In contrast, this paper’s focus on metric sampling uncertainty primarily pertains to model search and selection. The remainder of this paper presents the resolving power methodology, demonstrates its application to the AUROC and the AUPRC, and then reflects on the implications for model search.
2 ROC and PR curves
Our interest is in models that map cases to predicted classes. A discrete classifier is one that only outputs a class label. Applying a discrete classifier to test data produces a 2x2 confusion matrix (Table 1), with rows corresponding to the predicted class and columns giving the true class. A scoring classifier outputs a number on a continuous scale, such as an estimated probability, that represents the degree to which a case belongs to a class [6]. Applying a decision threshold to a scoring classifier produces a discrete classifier. Hand [10] shows that choosing a particular threshold is equivalent to specifying the relative costs of false positives versus false negatives.
Table 1: Example confusion matrix
actual + actual -
predicted + TP FP
predicted - FN TN
total P N
A variety of familiar evaluation metrics may be calculated for discrete classifiers such as accuracy, recall (hit rate, sensitivity, true positive rate), precision (positive predicted value), specificity, and the F1-score. These are known as single-threshold (or threshold-dependent) metrics. In contrast, threshold-free metrics use the full range of the original scores. Examples include the AUROC, the AUPRC, and the area under the precision-recall-gain curve (AUPRG), among others. Threshold-free metrics are advantageous since they allow users to adapt the model to a specific context [7]. The AUROC and AUPRC are preferred metrics when the primary goal is to achieve good discrimination so that cases are efficiently sorted into the positive and negative classes.
The ROC curve depicts the trade-off between the true positive rate (tpr) on the y-axis and the false positive rate (fpr) on the x-axis. A discrete classifier only gives a single point in ROC space, corresponding to its one confusion matrix. A scoring classifier gives points for every possible confusion matrix that can be formed by varying the decision threshold. The empirical ROC curve interpolates between these points to create a step function. As the number of points become arbitrarily large the empirical curve will approach the population ROC curve.
If a decision threshold is selected to flag 50 percent of all cases and the classifier is no better than random guessing then we expect it to identify half of the positives and half of the negatives, yielding the point $(0.5,0.5)$ in ROC space. Similarly, a random classifier flagging 20 percent of cases is expected to have a recall of 20 percent and a false positive rate of 20 percent. The random guessing classifier, then, is given by the $y=x$ line in ROC space. A perfect classifier ranks all positive cases above all negative cases, so it corresponds to the step function from $(0,0)$ to $(0,1)$ for all the positives, and then from $(0,1)$ to $(1,1)$ for all the negatives. Classifiers that lie above the identity line but below the perfect step function represent intermediate performance with better classifiers containing points closer to the $(0,1)$ northwest corner of ROC space.
The AUROC summarizes a classifier’s performance across all decision thresholds and is found by integrating the ROC curve over the $[0,1]$ range of false positive rates. Larger AUROC values are better, with the random classifier giving an $\text{AUROC}=0.5$ and the perfect classifier giving an $\text{AUROC}=1$. The AUROC, as a scalar value, is especially relevant for model tuning and selection. A disadvantage of the AUROC is that it can conceal local differences in performance. For instance, one classifier may be better for highly ranked cases while another is better for those in the intermediate or lower ranks. An important statistical property of the AUROC is that its value equals the probability that a classifier will rank a randomly chosen positive case higher than a randomly chosen negative case [9, 12].
The AUROC can be interpreted as an average sensitivity, assuming all specificity values are equally likely [10]. Several authors argue that the AUROC is deficient in key respects. swamidass2010croc remark that treating all specificity values as uniformly important is not appropriate for most problems and propose the area under the concentrated ROC as a corrective. Byrne [3] shows that the AUROC is typically not a proper scoring function. The AUROC has also been criticized as incoherent since two classifiers with the same AUROC will typically imply different relative costs of misclassification [10, @hand2023notes]. Hand proposes the H-measure as a coherent alternative (Online Resource 1 applies resolving power to the H-measure). In response to Hand, ferri2011coherent argue for alternative interpretations of the AUROC that are both coherent and model independent.
Precision-recall (PR) graphs plot precision on the y-axis and recall on the x-axis. In PR space a random classifier corresponds to the horizontal line $y=\frac{P}{P+N}=\text{prevalence}$ where $P$ is the number of positive cases and $N$ is the number of negative cases. PR curves are sensitive to class skew (meaning one class occurs more than the other) while ROC curves are not. This is because inputs to the ROC curve, the true and false positive rates, only depend on the column sums of the confusion matrix. Precision depends on the row sum of true and false positives, so all else equal, it will decrease with decreasing prevalence. Insensitivity to skew has been described as both an advantage [6] and disadvantage [17] of the ROC curve.
Just like the AUROC, the AUPRC reduces a scoring classifier’s performance to a single value, with larger values indicating better performance. Similar to the AUROC, the AUPRC can be interpreted as the classifier’s average precision over the $[0,1]$ range of recall values. Davis and Goadrich [4] demonstrate that there is a one-to-one correspondence between empirical ROC and PR curves since they both chart a unique mapping from confusion matrices to points in ROC or PR space. They go on to show that the AUROC and AUPRC give the same model rankings when one model’s curve “dominates” another’s. Informally, one curve dominates another if it lies above or equal to it across their domains. A dominating ROC curve will be northwest of the dominated curve, where its tpr is higher, its fpr is lower, or both. And a dominating PR curve will be northeast of a dominated curve, with higher precision, recall, or both across the entire domain. When there is no domination (when two curves cross) the AUROC and AUPRC can give different rankings. In cases of disagreement, the AUPRC favors classifiers with better performance in the early retrieval area, which is the region of low false positive rates in ROC space.
Because it gives more weight to the early retrieval area, the precision-recall curve is often recommended for highly-skewed datasets. Yet the empirical PRC is an imprecise estimate of the true curve, especially for small sample sizes and with strong class imbalance [2]. This raises the question of whether the advantages of the PRC are worth its cost in precision. Answering this question requires that we compare metrics measured on different scales.
3 Mapping between metrics
ROC analysis was initially developed to evaluate electronic sensors, such as radar, during World War II. In the 1950s research psychologists elaborated ROC analysis under the rubric of signal detection theory (SDT), which soon became influential within experimental psychology, psychophysics, and cognitive neuroscience [20]. Fundamental to SDT is the specification of two probability distributions: A noise distribution for trials when the signal is absent and a signal distribution for trials when the signal is present [9]. The binormal model (two Gaussians) is the most common choice for the signal and noise distributions. The SDT framework can be described in the language of binary classification with signal and noise trials considered members of the positive and negative classes, respectively.
A classification model applied to feature measurements generates class score distributions. For a simple example, suppose the two classes are women and men and that there is one feature of height. The classification model will just be the identity mapping applied to the height measurements. The binormal model should then be a good approximation for the score distributions. [^2] Height is believed to result from the sum of a large number of independent genetic and environmental effects, so by The Central Limit Theorem the distributions should be approximately normal. Figure 1 shows the binormal model for this example, using height distribution parameters from Our World in Data [16]. Women have an average height of 164.7 cm with a standard deviation of 7.1 cm, while men have an average height and standard deviation of 178.4 cm and 7.6 cm, respectively. The vertical dashed line is an example of a decision threshold, where any person above 171 cm is classified as a man and any below as a woman (this type of rule might be used in low visibility contexts where height is the most salient feature). Hit rates and false alarm rates can be calculated for that decision threshold, giving one point in ROC space.
Figure 1: A binormal classifier example. The distribution of men’s and women’s heights approximately follow a normal distribution. The model implies an AUROC of .906. The vertical dashed line at 171 cm is an example decision threshold.
Now we must address what we mean by “model quality”. This paper adopts a discriminative conception: Better classifiers yield greater separation in the class scores [11]. Importantly, model quality refers to the out-of-sample class score distributions, which are typically estimated with resampling methods or by using a test set. A perfect classifier completely separates the class scores while a random classifier has identical class score distributions.333hand2023notes formally define a random classifier as one with identical cumulative distribution functions (cdfs) so that $F_{0}\equiv F_{1}$, where $F_{0}$ and $F_{1}$ are the cdfs of the negative and positive class scores, respectively. We can also use cdfs to define a perfect classifier: One where there exists a decision threshold $t^{*}$ such that $F_{0}(t^{*})=1$ and $F_{1}(t^{*})=0$. Finally, if the cdfs are not identical and they do not completely separate the class scores then they partially separate the distributions. That is, there is partial separation if $F_{0}\not\equiv F_{1}$ and there is at least one threshold $t^{\prime}$ such that $F_{0}(t^{\prime})<1$ and $F_{1}(t^{\prime})>0$. Classifiers that are neither random nor perfect are those that partially separate the score distributions.
[…]
The mapping between metrics depends on how the score distributions are separated and we cannot know in advance how classifier improvement will change the risk scores. There are myriad ways to increase class separation, each indicative of different types of improvement. This paper solves this ambiguity by fiat: It assumes that simple manipulations of score distributions are a reasonable description of model improvement. A simple approach, used below, is to add fixed increments to the positive class scores. A concern is that a simple additive mechanism may poorly describe how improvement happens in practice. A more realistic, though more involved strategy is to base the mechanism on the observed early stages of model improvement (Online Resource 1 sketches this approach).
More generally, resolving power depends on two key assumptions: 1. That model improvement is a mostly homogeneous process, and 2. That we can approximate this process. The first assumption holds that improvement largely occurs in similar ways across different algorithms or hyperparameter settings. Though it does allow for random deviations due to sampling or to the intrinsic stochasticity of some algorithms. The second assumption is that we can do a decent job describing the common improvement process by finding the right sequence of class score distributions.
We have identified the quality dimension as an ordered sequence of distributions, but how should we measure location on this dimension? One option is to just use the model rankings themselves, which forms an ordinal scale [18]. Another option is to measure distribution overlap directly using the Bhattacharyya coefficient. Or we can use a measure that relates overlap to model quality, the AUROC and AUPRC being two examples among many. The AUROC has several properties that make it a good choice. Since the AUROC is an area (and a probability), equal differences across the scale represent equal differences in amount. Another advantage, mentioned above, is that it is unaffected by outcome prevalence. The AUROC is also the most popular threshold-free evaluation metric for binary classifiers, making it a natural choice for the reference metric.
For our purposes, the AUROC’s biggest advantage is that it is agnostic with respect to where changes occur in the score distributions. This fact is easiest to demonstrate with an empirical score distribution, defined as a finite set of risk scores and associated outcomes. Briefly, suppose there are $n^{+}$ positive cases, $n^{-}$ negative cases, and that all risk scores are unique. Further suppose that the classifier is not perfect, so $0.5\leq\text{AUROC}<1$, and we want to improve this by perturbing the risk scores. If we sort all cases together into a single list ranked by score, then the smallest improvements occur by finding pairs of adjacent scores that are “out-of-order”, such that the negative case has a higher score than the positive case, and re-ordering these pairs. Re-ordering a single pair will improve the AUROC by $\frac{1}{n^{+}}\times\frac{1}{n^{-}}$ regardless of where the improvement occurs. This follows from the probabilistic interpretation of the AUROC: There are $n^{+}\times n^{-}$ unique ordered pairs of positive and negative scores, which forms the number of events in the sample space. So resolving one out-of-order pair increases the probability by $\frac{1}{n^{+}\times n^{-}}$. In contrast, the AUPRC will improve more for resolving out-of-order pairs that are among the highest-ranked risk scores.
To summarize this section’s key points: Classifier quality is gauged by its outputs, the class score distributions. A sequence of increasingly separated class distributions forms a common quality dimension that charts the relationship between different evaluation metrics. A key caveat is that the mapping between metrics is contingent on how the score distributions are separated. Several characteristics of the AUROC make it a good choice as the reference measure of model quality. [^4] Note that these advantages pertain narrowly to the AUROC’s role as reference measure and do not counter to the criticisms of the AUROC referenced in the previous section. In particular, the AUROC always improves by a constant amount when resolving a pair of adjacent out-of-order risk scores. Moreover, the viability of resolving power does not hinge on using the AUROC as the reference. Instead, it is the sequence of class score distributions that are fundamental. Other metrics can be substituted so long as they form a mapping from the class score distributions to an interval scale.
4 Resolving power
The resolving power method can be summarized in four steps:
  1. 1.
    Sampling model: Specify class score distributions, prevalence, and sample size.
  2. 2.
    Signal curves: Use the sampling model to create a fine grid of improving classifiers. Find each metric’s values across the grid.
  3. 3.
    Noise distributions: Estimate metric sampling uncertainty by drawing random samples at points of interest within the quality grid.
  4. 4.
    Comparison: Use the above results to estimate and compare resolving power.
This section illustrates the core mechanics of the approach using an idealized example while later sections move to the applications. Suppose we are interested in comparing the resolving power of the AUROC with that of the fictional “area under the super-great curve” (the AUSGC). We construct a sampling model by specifying the class score distributions, prevalence, and sample size. Next, we create a grid of 1000 models. The first model has identical class score distributions for the random classifier and then we gradually shift the distributions apart so that the 1000th model has almost no overlap. Finally, we want to assess sampling models that give AUROC values of 0.7 and 0.9. We draw many replicates from these two models to estimate the sampling variability of the two metrics. Figure 2 summarizes the analysis.
Figure 2: Signal curve example. The two panels are united by the same sequence of models used to construct the quality grid. The AUROC serves as the common reference scale on the x-axis.
Following the previous section’s recommendation, the AUROC on the x-axis serves as the reference scale for model quality. The left panel is then just the identity mapping. The right panel shows how the AUSGC changes relative to the AUROC, giving the relative signal of the two metrics across the quality continuum. Unit slope indicates equal signal, a slope less than 1 favors the AUROC, and a slope greater than 1 favors the AUSGC. For a given point on the curve, repeated draws from the sampling model estimate each metric’s noise distribution. The signal curves allow us to map each metric’s uncertainty interval to a model quality interval, which forms the common basis for comparison.
[…]
A disadvantage of the resolving power definition is that it requires choosing an arbitrary $\alpha$ level. Appendix A describes an alternative approach that eliminates this requirement by expressing resolving power as a scaled standard error. This comes at a cost of stronger assumptions: The alternate approach assumes that the signal curve is locally well-approximated by a straight line and that the evaluation metric’s sampling distribution is roughly symmetric.
Returning to the example, for the AUROC 0.7 model shown in blue we have:
  • An AUROC of .7 with a 95% confidence interval of [.65, .75].
  • An AUSGC of .063 with a 95% confidence interval of [.013, .114]. This maps to an AUROC interval of [.53, .76].
[…]
  • An AUROC of .9 with a 95% confidence interval of [.85, .95]
  • An AUSGC of .4 with a 95% confidence interval of [.35, .45]. This maps to an AUROC interval of [.89, .91].
[…]
5 Binormal model
The binormal model, as the most commonly used in ROC analysis, serves as a good initial application of the approach. All code and data used below are available on GitHub. [^5] https://github.com/colinbeam/resolving_power Now we apply the four steps of the resolving power method.
Step 1: The sampling model. Assume a binormal model where negative class scores have a standard normal $\mathcal{N}(0,1)$ distribution and positive class scores have $\mathcal{N}(\delta_{i},1)$ distributions. The analysis explores a range of prevalence comprising the values [.01, .05., .10, .20, .30, .40, .50], which is the same set used by Mazzanti [14]. We explore a moderately sized classification task of 10,000 instances, so the lowest prevalence condition has 100 instances in the positive class.
[…]
$\delta_{i}=\sqrt{2}\times\Phi^{-1}(\text{AUROC}_{i})$
where $\Phi^{-1}$ is the inverse cumulative standard normal distribution. Note that an evenly spaced AUROC grid will require progressively larger shifts between class distributions as model quality increases. Next, we need to find the AUPRC values associated with each AUROC grid point. The AUPRC can be found from a binormal model via numerical approximation. Let $\alpha$ represent the outcome prevalence and $\Phi_{+}$ and $\Phi_{-}$ represent the cumulative Gaussian distributions for the positive and negative classes, respectively. Brodersen et al [2] derive the PR curve by finding precision (PPV) as a function of recall (TPR):
$\text{PPV}=\frac{\alpha\text{TPR}}{\alpha\text{TPR}+(1-\alpha)\left(1-\Phi_{-}\left(\Phi_{+}^{-1}(1-\text{TPR})\right)\right)}$
And to find the AUPRC they numerically approximate the integral:
[…]
To summarize the steps: First we create an evenly spaced grid of AUROC values using the implied shift parameter values from equation (1). We then use the shift values in equation (2), specifically for the $\Phi_{+}$ parameterization. This gives us the PR curve so that we may use equation (3) to find the associated AUPRC value.
Figure 3 shows the binormal signal curves for each condition. The relationship between metrics becomes more curvilinear as prevalence decreases. This implies that, all else equal, the AUPRC will be relatively more discriminating among higher quality models applied within low prevalence contexts. The signal curve for a prevalence of .5 is approximately a straight line with an intercept of zero and a slope of one—the identity mapping. Thus, the AUPRC and AUROC are estimating the same quantity but by using different formulas. In this condition, then, differences in resolving power will be due to differences from sampling error alone.
Figure 3: Mapping between AUROC and AUPRC for the binormal model.
Step 3: Noise distributions. We wish to assess a range of quality values by evaluating models with AUROCs of [.65,.75,.85,.95].65.75.85.95[.65,.75,.85,.95][ .65 , .75 , .85 , .95 ]. In the figures below these models are respectively labeled “Poor”, “Fair”, “Good”, and “Excellent”. For the four points of model quality we take 10,000 random samples from each implied binormal model and estimate AUROC and AUPRC values with the PRROC R package [8]. The AUPRC is estimated using the Davis and Goadrich method [4]. Ninety-five percent confidence intervals are found from the .025 and .975 quantile values of the simulation samples.
Step 4: Comparison. For the final step we use the curves in Figure 3 to map the AUPRC 95 percent confidence interval to the AUROC scale. We then find the relative difference in metric resolution with the AUROC as the baseline (equal to the relative difference in resolving power with the AUPRC as baseline):
$\Delta=\frac{\kappa_{\text{PRC}}-\kappa_{\text{ROC}}}{\kappa_{\text{ROC}}}=\frac{1/\kappa_{\text{ROC}}-1/\kappa_{\text{PRC}}}{1/\kappa_{\text{PRC}}}$
The simulation was repeated three times and estimates were averaged to smooth out their variability across runs.
Figure 4: Relative metric resolution by outcome prevalence and model quality for a binormal model with a sample size of N = 10,000. At each level of model quality 10,000 simulations are taken from the sampling model. Confidence limits are the .025 and .975 quantile values of the simulation samples.
Simulation results are shown in Figure 4. Beginning with the prevalence = .5 “identity mapping” condition, we see that the AUPRC is usually around 10 percent more variable than the AUROC, though the disadvantage is smaller in the “Excellent” model condition. In the remaining conditions the AUPRC suffers a greater disadvantage in the flatter portions of the signal curves, corresponding to contexts of low prevalence and poor model quality. Specifically, the AUPRC is at a disadvantage for all poor (AUROC = .65), fair (AUROC = .75), and good (AUROC = .85) models across all levels of prevalence. AUPRC resolution is typically about 10 percent larger, though in the flattest portion of the signal curve—the low prevalence and low model quality condition—the disadvantage is around 30 percent. The AUPRC has better resolving power only for excellent models (AUROC = .95) applied to moderately to strongly skewed datasets (prevalence of .2 and below).
Figure 4 shows relative performance, but it is also important to consider how absolute uncertainty varies across conditions. Figure 5 explores these relationships using Hanley and McNeil [12]’s formula for the approximate standard error of the AUROC. The standard error is strongly decreasing in prevalence, shown by the vertical separation between lines. The .01 condition has an especially large standard error, making relative imprecision even more costly in absolute terms. The standard error is mostly decreasing in model quality, though interestingly, slightly increases from an AUROC of 0.5 before reaching a maximum around 0.6. As an aside, we can also form “normal approximation” confidence intervals by taking plus or minus 1.96 times the standard errors from Figure 5. The normal approximation intervals are typically close to the simulation confidence intervals. For a prevalence = .01 and AUROC = .65 the simulation 95 percent confidence interval is [0.596, 0.702] while the normal approximation is [0.591, 0.709]. The approximation becomes worse as the AUROC increases: In the prevalence = .01 and AUROC = .95 condition the simulation confidence interval is [0.929, 0.967] while the normal approximation is [0.92, 0.98]. The adequacy of this approximation bears on the utility of the alternative method for estimating resolving power, described in Appendix A.
Figure 5: The relationship between AUROC and its standard error for different levels of outcome prevalence. The standard error is found using Hanley and McNeil (1982)’s approximation formula.
Thus, for moderately sized (N = 10,000) classification tasks the AUROC will typically provide better resolution. Importantly, these results are essentially unchanged for different sample sizes. For both a magnitude smaller (N = 1000) and larger (N = 100,000), the AUROC is generally better, with the AUPRC showing a relative advantage only among excellent models with an outcome prevalence of 20 percent of less. Appendix B presents results for these additional scenarios.
Now suppose the AUPRC is what we really want to maximize. That is, the AUPRC best captures our intuitions about what makes a good model. But also suppose that we know our resolving power assumptions are true: That risk scores follow a binormal distribution and model improvements come from additive shifts. Then what this section’s results tell us is that we should use the AUROC for search even if our goal is to find the best AUPRC model (at least in most contexts). Under this hypothetical scenario we don’t need to balance the AUROC’s greater resolving power against our preference for the AUPRC because we know precisely when the AUROC is a better guide. But once we select the final model we should still describe its performance with the AUPRC since that is what we really care about.
Moving from the hypothetical to reality, we know that this section’s assumptions will never fully hold. This means that the results should be taken as only general guidance to be weighed against other criteria. For instance, the AUPRC suffers only a modest disadvantage for “good” models with moderate prevalence. Once we factor the influence of assumptions violations, which we could test with sensitivity analyses that explore other paths towards class score separation, the disadvantage may disappear.
It is uncertain how robust this section’s guidance is to deviations from the binormal model. One way to address this concern is to replace the binormal with a domain-specific data-generating process. However, most applications will not have the requisite quantitative framework to accomplish this. An alternative strategy is to start with a dataset and a baseline model and then build the sampling model from an initial set of risk scores. This empirically-driven approach is explored in the next section.
6 Empirical sampling models
This section shows how specific problem information can be incorporated into the resolving power approach. We will explore an example task where the aim is to predict 30-day hospital readmissions among diabetes patients using features such as patient demographics, prior utilization, diagnoses, lab tests, and medications. The data for this example can be found at the UCI Machine Learning Repository. [^6] For access and a description of the original dataset go to https://archive.ics.uci.edu/ml/datasets/diabetes+130-us+hospitals+for+years+1999-2008. The post-processed data can be found at the GitHub address listed above. After applying recommended restrictions, the dataset includes 69,973 total records with 6,277 readmissions for an outcome prevalence of about 9 percent. So this is an example of an imbalanced class problem for which the AUPRC is often recommended.
Now how can we use the data to guide our choice of a sampling model? A seemingly sensible approach is to fit an initial classifier and then use its risk scores to inform the choice. The initial model might be the simplest algorithm among a set of candidates, or it could be a preferred algorithm using its default hyperparameters. We may then construct a sampling model from the empirical distribution (the set of outcomes and risk scores) in a couple of different ways. One is to find a parametric model that gives a good approximation to the empirical distribution. Another is to treat the empirical distribution as the population, as is done in resampling methods such as bootstrapping.
For the readmissions data, we use a simple logistic regression as the initial model, estimating risk scores using 5-fold stratified cross-validation. The estimated AUROC is .646 for this initial model. By the previous section’s taxonomy, this is a “poor” model with about a 10 percent outcome prevalence. So the binormal model results suggests that we should prefer the AUROC to the AUPRC.
Figure 6: Top panel: Estimated logit effects with kernel density estimates for the positive and negative class. Bottom panel: Q-Q norm plots of sample versus theoretical quantiles. Straight lines pass through the 1st and 3rd quartiles.
The distribution of patient effects are shown in Figure 6, with a rug plot and density estimates in the top panel and qqnorm plots below. A normal approximation does not appear appropriate as both the positive and negative class have large clusters of scores in the lower tail. We could hunt for a better parametric approximation—perhaps some type of mixture distribution—but instead we will use the empirical distribution as the population sampling model.
Moving to the second step, how can we use an empirical sampling model to construct a signal curve? A simple option is to add small increments to all positive class scores, which is analogous to the approach we used with the binormal model. Incremental improvement will then be concentrated among the negative and positive cases with the closest risk scores, wherever they may reside in the distribution. This seems reasonable as it assumes that marginally better models will first amend the ranking of cases that need the smallest adjustments.
The previous section created an evenly-spaced grid of AUROC values using binormal model analytic results. An evenly-spaced grid is also possible for an empirical distribution: To begin, suppose there are $n^{+}$ positive cases with risk scores $r_{i}^{+}$ for $i\in\{1,...,n^{+}\}$ and $n^{-}$ negative cases with risk scores $r_{j}^{-}$ for $j\in\{1,...,n^{-}\}$. Further suppose all risk scores are unique and that the classifier is not perfect, so $r_{i}^{+}<r_{j}^{-}$ for at least one $(i,j)$ pair. This implies that $0.5\leq\text{AUROC}<1$ for the initial AUROC value. We will build the grid in the direction of improving AUROC, though it is straightforward to adapt the process for decreasing AUROC. First, find $\delta_{1}=\text{min}\left(r^{-}_{j}-r^{+}_{i}|r^{-}_{j}>r^{+}_{i}\right)$, so $\delta_{1}$ is the smallest positive difference in risk scores between two cases that are out-of-order such that the negative case is assigned higher risk than the positive case. Similarly, we can find $\delta_{2}$ as the second smallest difference, $\delta_{3}$ as the third smallest, etc. Now if we add $\delta_{1}+\epsilon$ to all positive class risk scores where $\delta_{1}<\delta_{1}+\epsilon<\delta_{2}$ we will shift the positive distribution just enough to resolve one pair of out-of-order risk scores, but no more. From above we know that the AUROC will then increase by $\frac{1}{n^{+}}\times\frac{1}{n^{-}}$. If instead we had added $\delta_{2}+\epsilon$ with $\delta_{2}<\delta_{2}+\epsilon<\delta_{3}$ then it would have fixed two pairs of scores and the improvement would have been $\frac{2}{n^{+}\times n^{-}}$. Thus, we can precisely increase or decrease the AUROC in $\frac{1}{n^{+}\times{n^{-}}}$ increments. The result is useful for determining an initial increment to shift the class scores. Achieving a fixed increment across the grid requires updating the score distance calculations after each step, but this is computationally costly and is typically unnecessary. Instead, it is most important to choose an initial increment that creates a high density of points across the grid range so that the signal curve may be reliably estimated.
Figure 7 shows the signal curve constructed from shifting the readmissions class score distributions. The starting model AUROC and AUPRC values of .646 and 0.166 are shown by the dot. The curve is built by shifting the positive class distribution above and below the starting point, using an initial increment that produces a change of .001 AUROC units. Each empirical distribution along the grid is considered the population, so the associated population AUROC and AUPRC are just the sample values. There are a total of 1000 grid points, which range from .54 to .92 in AUROC and from .12 to .50 in AUPRC. In practice, it is rare to see substantial improvement from initial performance, so these ranges cover a larger space than is expected to be observed during model search. The shape of the curve in Figure 7 is similar to the binormal signal curves: For lower AUROC values the slope is relatively flat but then it increases with improving model quality.
Figure 7: An empirical signal curve from the readmissions data logistic regression model. The black point shows baseline performance. The curve is constructed by incrementing positive scores above and below the baseline distribution.
Moving to the third step of noise estimation, the initial model is the natural choice to evaluate since improvements will be made from this starting point. We generate 10,000 samples from the empirical distribution, fixing the prevalence with stratified sampling, and then find 95 percent confidence intervals. The fourth and final step uses the signal curve to estimate and compare metric resolution. Results of the sampling experiment are displayed in Table 2.
Table 2: Simulation results summary. Lower and upper CI bounds are for 95 percent confidence intervals.
metric Lower CI Upper CI $\kappa$ resolving power
AUROC 0.6391 0.6535 0.0144 69.4
AUPRC 0.1601 0.1732 NA NA
AUPRC to AUROC 0.6341 0.6591 0.0250 40.0
The last row uses the signal curve to map AUPRC to the AUROC scale. Recall that metric resolution, $\kappa$, is just the width of the 95 percent confidence interval in AUROC units. The AUROC has considerably better discrimination with a resolving power that is over 70 percent greater than the AUPRC. The absolute difference in confidence interval widths is about 0.011 AUROC units, which could be substantial relative to the often small improvements achieved during hyperparameter tuning. Hence, the resolving power analysis predicts that the AUROC will be better than the AUPRC for model search. But is it actually? The next section considers the challenge of empirically assessing the resolving power framework.
7 Empirical validation
Resolving power’s empirical importance crucially depends on the topography of the model search space. Its impact will be limited when search occurs in either “signal dominant” or “noise dominant” contexts. For instance, if the model evaluation points are spread across a steep region of the space then the model quality signal will overwhelm the noise for all metrics. In this case there will be consensus among metrics so it will not matter which is used for model selection. Conversely, some evaluation points may span an optimum that forms a flat region of the search space. If the flat portion is the entire search space or is sufficiently elevated (imagine an isolated butte) then again the choice of evaluation metric will not matter. But instead of unanimity we should now expect the metrics to agree only some of the time. [^7] If the metric estimates are independent across the optimum then the probability of consensus will be $(1/n)^{k-1}$ where $n$ is the number of evaluation points on the optimum and $k$ is the number of metrics.
The above concerns noted, we still want to test the hypothesis that higher resolving power leads to better model selection. An immediate challenge is that for actual data, such as with the diabetes readmissions example, we do not have access to the population. Yet with a sufficiently large dataset we can approximate drawing samples from a known population. This is the approach kohavi1995study uses to compare how different resampling methods affect classifier accuracy. Since the population characteristics of real data cannot be known, Kohavi uses the holdout method to estimate the “true” accuracies with relatively large test sets.
[…]
The AUROC had greater baseline estimated resolving power than the AUPRC on 499 of the 500 trials, meaning that the AUROC would be preferred for model search. For simplicity, the one aberrant trial is excluded from the subsequent analysis. The two metrics agreed on the best model on 281 out of the remaining 499 trials. Of the 218 disagreements, using the AUROC for model selection led to better AUROC test set performance on 146 of the trials and better AUPRC performance on 134 of the trials. That is, it was better on 67 and 61 percent of trials, higher than the predicted 50 percent if the two metrics were equally good for search.
Table 3: The top row gives the average performance for trials where ROC and PRC model choice disagrees. Disagreement occurred on 218 out of 499 total trials. The bottom row gives performance on the trials where the two metrics agree.
agree tune ROC, test ROC tune PRC, test ROC tune ROC, test PRC tune PRC, test PRC
FALSE 0.6205 0.6168 0.1495 0.1478
TRUE 0.6227 0.6227 0.1509 0.1509
[…]
For a given application the importance of resolving power will depend on the classification task, the candidate evaluation metrics, and the set of models being compared. Consequently, the observed effects of metric choice, such as those from this section, are liable to shrink or grow with any changes made to the model search space. This sensitivity to the problem context poses a special challenge for assessing the general empirical importance of resolving power.
8 Conclusion
Evaluation metrics form the contours of a model’s performance topography, so choosing the right metric is essential for successful navigation of this space. Resolving power is a framework for comparing threshold-free metrics. Central to the method is the specification of a class score sampling model that is used to both manipulate model quality and probe sampling variability. The quality dimension, which serves as the standard for comparison, is an ordered sequence of increasingly separated score distributions. A pivotal assumption is that movement on the quality dimension resembles the actual process of model improvement. This paper uses simple additive shifts to separate the class scores. Future work can test and refine this assumption by observing how risk scores evolve as algorithms learn from real-world data. Signal curves show how evaluation metrics respond to changes in classifier quality. Metric error variance is found with random draws from the sampling model. Resolving power is classifier-agnostic since it operates on risk scores that are downstream of a classification model. Binormal model simulation results provide general rules-of-thumb for when the AUROC will have stronger resolving power than the AUPRC. The empirical method allows researchers to use their data and an initial classifier of their choice to construct a sampling model for estimating resolving power.
If we imagine evaluation metrics occupying a 2-dimensional space with interpretability measured on the x-axis and resolving power on the y-axis, then a metric’s (intepretability, resolving power) coordinates describe its usefulness for model search. A (low, low) metric has no value. A (low, high) metric reliably maximizes an irrelevant or inscrutable target. A (high, low) metric fails to maximize the desired target. Of course (high, high) is best, but the challenge remains on how to create these types of metrics.
[…]
Supplementary information
All code and simulation data used in the paper are available on GitHub at https://github.com/colinbeam/resolving_power
Appendix A Linear approximation method for resolving power
The resolving power approach from the main text has a couple of disadvantages: It requires estimating the signal curves over many grid points and choosing a specific $\alpha$ value for the confidence interval width. This section outlines a local, linear approximation that avoids both of those drawbacks. Return to the toy example from Section 4, where we wish to compare the resolving power of the AUSGC versus the AUROC. The linear approach is easier to express by flipping the axes, plotting the AUSGC on the x-axis and the AUROC on the y-axis, as shown in Figure 8.
Suppose we want to compare the resolving power of the AUSGC versus the AUROC at the point (0.16, 0.8) shown in black. Figure 8 shows the full signal curve, but we only need to evaluate a few points to estimate the tangent line shown in red. The slope at the evaluation point, which gives the relative signal of the two metrics, is 0.69. Since the slope is less than 1 it means the AUSGC has a relatively stronger response to improvements in model quality.
Next, we must estimate sampling variability at the evaluation point. Again, suppose we form many simulation samples of the evaluation metrics. But instead of finding percentile confidence intervals, we use the simulation samples to estimate the standard deviation of the AUSGC and AUROC, denoted respectively as $\hat{\sigma}_{\text{S}}$ and $\hat{\sigma}_{\text{R}}$. [^9] The standard deviation of the simulation samples estimates the standard error of the evaluation metric. If we assume that the distribution of the sample AUROC is approximately normal, then we can form a $1-\alpha$ confidence interval by selecting a $z_{(1-\alpha/2)}$ critical value and multiplying the standard error. Using $z_{(1-\alpha/2)}=1.96\approx 2$ gives an approximate 95 percent confidence interval. So for the AUROC we get the interval: $[-2\hat{\sigma}_{\text{R}},\;2\hat{\sigma}_{\text{R}}]$. The metric resolution (the width of the confidence interval in AUROC units) is then $\kappa_{\text{ROC}}=4\hat{\sigma}_{\text{R}}$. Similarly, the approximate 95 percent confidence interval for the AUSGC is $[-2\hat{\sigma}_{\text{S}},\;2\hat{\sigma}_{\text{S}}]$. Now we use the linear approximation to map AUSGC to the AUROC scale. Suppose the slope of the linear approximation is $\beta_{1}$ and the intercept is $\beta_{0}$, then we obtain the confidence interval $[-2\hat{\sigma}_{\text{S}}\beta_{1}+\beta_{0},\;2\hat{\sigma}_{\text{S}}\beta_{1}+\beta_{0}]$ and its width is $\kappa_{\text{SGC}}=4\hat{\sigma}_{\text{S}}\beta_{1}$. Taking the ratio of metric resolutions gives:
$\frac{\kappa_{\text{SGC}}}{\kappa_{\text{ROC}}}=\frac{4\hat{\sigma}_{\text{S}}\beta_{1}}{4\hat{\sigma}_{\text{R}}}=\frac{\beta_{1}\hat{\sigma}_{\text{S}}}{\hat{\sigma}_{\text{R}}}$
The $z$ critical value cancels and we are left with the ratio of standard errors scaled in AUROC units. Thus, the linear approach requires only comparing the ratio of the standard errors to the slope of the signal curve, eliminating the need for an $\alpha$ level. Using the Leibniz notation $\beta_{1}=\frac{dR}{dS}$, we must only check the inequality:
$\frac{dR}{dS}>\frac{\hat{\sigma}_{\text{R}}}{\hat{\sigma}_{\text{S}}}$
Inequality A.1 makes transparent the signal to noise comparison: The left side is the relative signal of the two metrics while the right side is the relative noise. If the inequality holds it means that the signal of the AUROC overwhelms its noise, giving it relatively greater resolving power.
Figure 8: Signal curve plotting the AUROC reference scale on the y-axis. A linear approximation of the signal function at the evaluation point is shown in red.
The simplicity of the linear approach bears both its strengths and weaknesses. Estimating the tangent curve requires only a few points in the immediate vicinity of the evaluation point, obviating the need to find the full signal curve or choose a confidence interval width. The approach crucially assumes:
  1. i.
    The metric sampling distribution is symmetric.
  2. ii.
    A line is a good approximation of the signal curve over the region of interest.
Recall that we calculate metric resolution by using the signal curve to map confidence interval limits from one scale to another. A linear function will give a satisfactory approximation when these confidence limits are narrow. However, it will be poor for wide confidence intervals bracketing a curve segment that has a rapidly changing slope.
Appendix B Additional binormal results
Figure 9: Relative metric resolution by outcome prevalence and model quality for a binormal model with a sample size of top panel: N = 1000 and bottom panel: N = 100,000. At each level of model quality 10,000 simulations are taken from the sampling model. Confidence limits are the .025 and .975 quantile values of the simulation samples.
We extend the binormal investigation to sample sizes that are an order of magnitude smaller and larger than those in the main text, shown respectively in the top and bottom panels of Figure 9. For the $N=1000$ study the simulation was again repeated three times with estimates averaged across runs. In the $N=100,000$ study the simulation was conducted only once.
Interestingly, for $N=1000$ the direction of the differences across conditions are the same as found in the $N=10,000$ study. Only the relative magnitudes have changed. Specifically, the AUPRC has superior resolving power only for “excellent” models with an outcome prevalence of 20 percent or less. The primary difference in magnitudes are found in the one percent prevalence condition, which now shows a smaller disadvantage for the fair to good models, and a smaller advantage for the excellent models. Note that there are now only ten instances in the positive class for one percent prevalence.
Results from the increased order of magnitude $N=100,000$ study are also similar. There is one condition where the direction of the effect has flipped—the excellent model condition with a prevalence of 30 percent—though the relative difference is essentially zero. The other primary difference is that the AUPRC is now at the biggest disadvantage for the one percent prevalence “fair” model condition.
In summary, differences in sample size for the binormal model generally do not affect the direction of the effects, only their relative size. The AUPRC maintains an advantage only for excellent models with an outcome prevalence of 20 percent or less.